OpenStudy (anonymous):

For each of the given functions f(x), find the derivative (f^-1)'(c) at the given point c, first finding a=f^(-1)(c) f(x)=4x+7x^9; c=-11 a=?

OpenStudy (anonymous):

whatever you do, do not try to find the inverse of this function, that is not the point of the exercise

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

you need only \(f^{-1}(-11)\) which you can do in your head

OpenStudy (anonymous):

how does that help?

OpenStudy (anonymous):

how do you find the inverse?

OpenStudy (anonymous):

because once you have it, you know how to find \[\frac{d}{dx}f^{-1}(-11)\]

OpenStudy (anonymous):

you do ( i repeat, do not) find \(f^{-1}(x)\) but only \(f^{-1}(-11)\)

OpenStudy (anonymous):

no no a thousand times no, you do not find \(f^{-1}(x)\) as an equation it is unnecessary and too much work and might not even be possible

OpenStudy (anonymous):

\[f^{-1}(-11)=-1\] by your eyeballs, because \(f(-1)=-11\)

OpenStudy (anonymous):

im so lost.

OpenStudy (anonymous):

satellite, she's doing the derivative of the inverse, not just the inverse \[\frac{ df ^{-1} }{ dx } =\frac{ 1 }{\frac{ df }{ dx } } \]

OpenStudy (anonymous):

therefore, since \(\frac{d}{dx}f^{-1}(x)=\frac{1}{f'(f^{-1}(x))}\) and since you know \(f^{-1}(-11)=-1\) your job is only to take the derivative of \(f\), evaluate it at -1 and take the reciprocal , that is all

OpenStudy (anonymous):

@Algebraic! yes i know, it is the derivative of the inverse evaluated at a number

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

it is this \[\frac{d}{dx}f^{-1}(c)=\frac{1}{f'(f^{-1}(c))}\]

OpenStudy (anonymous):

so we need only one number, namely \(f^{-1}(-11)=-1\) to do this problem

OpenStudy (anonymous):

1/f'(-1)?

OpenStudy (anonymous):

and therefore all you need is the derivative of \(f\) itself

OpenStudy (anonymous):

4+63x^8

OpenStudy (anonymous):

@monroe17 no, it is the reciprocal of the derivative of the original function, evaluated at the inverse function

OpenStudy (anonymous):

ok so you have \(f'(x)=4+63x^8\) right? what you need is \(f'(-1)\)

OpenStudy (anonymous):

you can do it with your eyes, i get \(f'(-1)=4+63=67\)

OpenStudy (anonymous):

ah I see what @satellite73 is saying...

OpenStudy (anonymous):

and so your answer to \[\frac{d}{dx}f^{-1}(-11)\] is \[\frac{1}{67}\] that is all finito halas stick a fork in it

OpenStudy (anonymous):

yeah, he's right, I misread the c= thing...

OpenStudy (anonymous):

so can you simplify all that down into steps for me..?

OpenStudy (anonymous):

again steps are find \(f^{-1}(-11)=-1\) take the derivative of \(f\) get \(f'(x)=4+63x^8\) evaluate at \(x=-1\) get \(f'(-1)=67\) take the reciprocal to get the derivative of the inverse evaluated at \(-11\) get \[\frac{1}{67}\]

OpenStudy (anonymous):

written in one line it is \[\frac{d}{dx}f^{-1}(-11)=\frac{1}{f'(f^{-1}(-11))}=\frac{1}{f'(-1)}=\frac{1}{67}\]

OpenStudy (anonymous):

the only hard part is finding \(f^{-1}(-11)\) but since \(f(x)=4x+7x^9\) the solution to \[4x+7x^9=-11\] sort of jumps out at you

OpenStudy (anonymous):

actually if you look at the way the question was written, it asked first for \(f^{-1}(-11)\) which is the correct first step. it called it \(a\) you need to do that first

OpenStudy (anonymous):

so for... x^2-13x+61 c=25.. = x^2-13x+61=25 f^(-1)(25)?

OpenStudy (anonymous):

yeah you are trying to solve \(x^2-13x+61=25\)

OpenStudy (anonymous):

x=9,4?

OpenStudy (anonymous):

if you know it, then you know \(f^{-1}(25)\) although this is a quadratic function so it doesn't have an inverse unless you restrict the domain

OpenStudy (anonymous):

on the interval [6.5, inf)

OpenStudy (anonymous):

yeah see the problem? you are not sure which one, because the inverse in this case is not a well defined function the quadratic is not one to one

OpenStudy (anonymous):

oh well then it is one to one i suppose pick 9

OpenStudy (anonymous):

since 9 is the on in the interval now look how easy the problem becomes now that you have it the derivative is easy evaluating it at 9 is easy taking the reciprocal is real easy, you just write it

OpenStudy (anonymous):

1/5 yay!

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

at first it looks confusing, then it is simple

OpenStudy (anonymous):

do you know why it works?

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

\[f(f^{-1}(x))=x\] by definition

OpenStudy (anonymous):

take the derivative of both sides. left is easy right uses chain rule \[f'(f^{-1}(x)\times (f^{-1}(x))'=1\] solve for \((f^{-1}(x))'\) get \[(f^{-1}(x))'=\frac{1}{f'(f^{-1}(x))}\]

OpenStudy (anonymous):

well i got my left and right backwards, but that is the idea now you can use it to find the derivative of any inverse function, assuming you know the derivative of the original function

OpenStudy (anonymous):

okay, thanks so much!